Find the equation to the line tangent to y = 1/(x+3) where x = 2

Use the definition of derivative to find the slope, lim laws, and algebra
Answer #1

Solution: 1 Tangent line to y= =—at a point x = 2 x+3 Put x=2 into the equation 1 y= 2+3 Y= 1 In order to find the slope of t